OpenStudy (anonymous):

what ratio is equivalent to 4:13

OpenStudy (texaschic101):

ratios can be written as frctions 4:13 can be written as 4/13....so basically, you are looking for equivalent fractions. To find an equivalent fraction, you have to multiply numerator and denominator by the same number... example : 4/13 * 2/2 = 8/26...so 8/26 is euivalent to 4/13...or 8:26 is equivalent to 4:13. 4/13 * 3/3 = 12/39..so 12/39 is equivalent to 4/13..or 12:39 is equivalent to 4:13. whatever number you multiply 4/13 by..it has to be the same number ..you could multiply by 4/4 or 5/5 and get other equivalent fractions which can be written as ratios. any questions ?

OpenStudy (texaschic101):

typo on first line...ratios can be written as FRACTIONS